# Think with 3D: Geometric Imagination Grounded Spatial Reasoning from Limited Views

### v1

@@ Abstract
Though recent advances in vision–language models (VLMs) have achieved remarkable progress across a wide range of multimodal tasks, understanding 3D spatial relationships from limited views remains a significant challenge. Previous reasoning methods typically rely on pure text (e.g., topological cognitive maps) or on 2D visual cues. However, their limited representational capacity hinders performance in specific tasks that require 3D spatial imagination. To address this limitation, we propose 3DThinker, a framework that can effectively exploits the rich geometric information embedded within images while reasoning, like humans do. Our framework is the first to enable 3D mentaling during reasoning without any 3D prior input, and it does not rely on explicitly labeled 3D data for training. Specifically, our training consists of two stages. First, we perform supervised training to align the 3D latent generated by VLM while reasoning with that of a 3D foundation model (e.g., VGGT). Then, we optimize the entire reasoning trajectory solely based on outcome signals, thereby refining the underlying 3D mentaling. Extensive experiments across multiple benchmarks show that 3DThinker consistently outperforms strong baselines and offers a new perspective toward unifying 3D representations into multimodal reasoning. Our code will be available at https://github.com/zhangquanchen/3DThinker.
@@ 1 Introduction
Recent advances have attempted to enhance the spatial reasoning capabilities of VLMs [45, 39, 68, 34, 24, 40, 13]. As illustrated in Fig. 1, existing methods can be broadly divided into two categories. The first category performs reasoning with pure text [7, 40, 14, 45, 39] or 2D visual cues [17, 13, 71], whose representational capacity for complex spatial layouts is inherently limited. To mitigate this limitation, methods such as MindCube [80] train models to generate cognitive maps of 3D layouts; however, they rely on bird’s-eye-vie (BEV) annotations to construct these maps. Ego3D [24] further employs external models—GroundingDINO [38] for referring expression comprehension (REC) and DepthAnythingv2 [76] for depth estimation, to automatically generate cognitive maps Yet, constrained by the performance of these models, such methods often fail on low-resolution or uncurated images. The second category incorporates auxiliary modalities as additional inputs (e.g., point clouds, camera parameters) [9, 32]. However, these settings restrict the model’s applicability in real-world scenarios where only monocular images are available. Moreover, several recent methods invoke external encoder or tool-usage to obtain prior information (e.g., encoded 3D tokens [21], depth maps [40, 13, 5]). Importantly, these techniques do not constitute an intrinsic capability of the model and introduce additional inference overhead.
@@ 2.1 Multimodal Reasoning
Large language models (LLMs) have experienced rapid development, and demonstrated strong performance across a wide range of tasks [11, 92, 85, 35, 82, 70, 16]. Building on these advances, recent works have highlighted that in-context learning, including intermediate rationales, can significantly enhance the performance of LLMs [49, 64, 22, 63, 37, 93, 81]. Current reasoning methods can be categorized into three types: pure-text, visual, and latent reasoning.
Pure-text reasoning:  [10, 56, 29, 27, 53, 3] elicit textual step-by-step reasoning inspired by [25]. They typically rely on textual descriptions, which can limit the reasoning capabilities when dealing with visual evidences that cannot be adequately described using pure textual language. Visual reasoning: to solve the problem mentioned above, some methods integrate visual evidences directly into the reasoning trajectory, whether in multi-hop or continuous modes. Some intrinsic multi-hop methods [39, 51, 12, 62], first generate detailed visual cues within the model itself (e.g., bounding boxes, coordinates, or masks), and then the further reasoning is conducted based on these cues. Other extrinsic tool-usage methods [44, 54, 67, 91], enhance the ”think with image” capability by dynamically invoking external image tools. On the other hand, continuous methods like GRIT [20] and SIFThinker [13] generate continuous visual reasoning to enable iterative corrections during the single-step reasoning process. Latent reasoning: some studies have shown that incorporating intermediate hidden representations into LLMs can effectively enhance model capabilities [4, 77, 18, 55].  [26] replaces CoT tokens with continuous latent embeddings, allowing unconstrained reasoning in the latent space to tackle complex tasks. More recently, Mirage [79] and LVR [31] utilize special visual tokens alongside ordinary text during reasoning. They explore visual information within the model by implicitly supervising the generation of image latent, thereby enabling reasoning with 2D visual latent.
@@ 2.2 Spatial Understanding
Spatial understanding encompasses skills such as 3D imagination and spatial cognition, which are essential for perceiving and manipulating spatial relationships in both 2D and 3D environments [72, 86, 6, 43, 58, 84, 42, 87, 19]. Recently, much efforts have been dedicated to evaluating the spatial understanding ability of VLMs [24, 80, 48, 90, 41, 89, 30]. Additionally, several methods have been proposed to enhance spatial understanding. For example,  [40, 5, 17, 45, 46] equip LLM with additional multiview, depth or point cloud inputs, essentially serving as input enhancement. Furthermore, 3DRS [28] introduces a teacher model for 3D supervision to achieve explicit spatial representation alignment; however, this method requires input that includes the 3D coordinates corresponding to each pixel. Moreover, VLM-3R [21] employs implicit 3D tokens from a pre-trained model (e.g., CUT3R [60]) to achieve spatial awareness by incorporating prior information, necessitating inference with extensively 3D foundation model. Recently, methods like MindCube [80] and Ego3D-VLM [24] have facilitated spatial understanding by constructing textual cognitive maps.
@@ 3.1 Data Generation
Due to the fact that VLMs naturally only generate textual tokens, they require additional supervised training to learn how to produce interleaved reasoning patterns that incorporate 3D information. Therefore, we synthesize specific training corpora based on the 10K training data from MindCube dataset [80]. Given an image set from different views $\mathcal{I}=\{I_{1},I_{2},\ldots,I_{n}\}$, a question $Q$, and the ground truth response $R$, we employ a high-level model (i.e., GPT-4o) $M$ to complete the reasoning chain. Specifically, we prompt the model $M$ to generate step-by-step reasoning that contains placeholders (3D special tokens), where these tokens represent imagined 3D scenes in the mind. Denote the response $o$ as:
@@ 3.2 Supervision for 3D Grounded Reasoning
The overall training objective incorporates both 3D alignment and textual losses, thereby enabling the model to seamlessly incorporate 3D imagining into its textual reasoning process. Here, $\lambda_{3D}$ and $\lambda_{text}$ serve as hyperparameters that balances coefficients.
@@ 3.3 Reinforced Spatial Mentaling
After the supervised training, the model has begun to exhibit the ability of 3D mentaling during thinking process. To further optimize the 3D visual token in the reasoning process, we can extract the last layer hidden state of 3D special token $t_{3D}$ (i.e., <|latent_start|><|latent_pad|> ... <|latent_end|>) in each trajectory $o$ and perform optimization. Specifically, the projected features $F_{proj}^{RL}$ are computed based on Eq. 4 at step $t$, with VGGT features $F_{3D}$ serving as constraints during the RL stage. That is, the cosine similarity between the VGGT features and the projected features is calculated to serve as the reward $r_{\text{3D}}$.
@@ 4 Experiments
Hyper-parameters. For 3DThinker, in the stage 1, we set the MLP depth to 6, with the learning rate of 1e-4, latent size of 12, epoch of 10. In Eqn. 9, the hyper-parameters $\lambda_{3D},\lambda_{text}$ are uniformly set to 0.1 and 1, respectively. In the stage 2, we set the balancing coefficient of all three reward to 1, with the learning rate of 1e-5, the rollout number of 8. Additional details are provided in the Supp. Mat..
@@
$\textit{{3DThinker}-S1}_{\textnormal{Qwen2.5-3B}}$ 53.2 54.8 74.5 52.3 59.5 37.7 55.3
$\textit{{3DThinker}-S1+S2}_{\textnormal{Qwen2.5-3B}}$ 59.1 60.2 78.4 58.2 64.7 41.9 60.4
@@
$\textit{{3DThinker}-S1}_{\textnormal{Qwen2.5-7B}}$ 57.3 61.5 77.9 56.3 61.7 41.5 59.4
$\textit{{3DThinker}-S1+S2}_{\textnormal{Qwen2.5-7B}}$ 63.7 68.3 81.1 63.3 68.6 43.3 64.7
@@ 4.1 Benchmarking Generalist VLMs
In this section, we comprehensively investigate different training stages in 3DThinker across various generalist VLMs. We conduct experiments on MindCube-Tiny [80] and Ego3D-Bench [24] benchmarks, both of which are designed to evaluate the spatial understanding ability from limited views.
@@ 4.3 Training Strategies
To further demonstrate the effectiveness of our training paradigm, we compare 3DThinker against several representative training strategies. Among them, Aug-CGMap-FFR-Out and Plain-CGMap-FFR-Out serve as SOTA baselines introduced in [80]. Specifically, Aug-CGMap-FFR-Out performs reasoning with the augmented cognitive map (camera-view information included), whereas Aug-CGMap-FFR-Out relies solely on plain cognitive maps without augmentation.
@@ 4.4 Visualization
We visualize the results of 3D mentaling in Fig. 4. During inference, we extract the last layer hidden states corresponding to the 3D special tokens. These 3D latents are projected into the VGGT feature space via the projector illustrated in Fig. 3. The projected features are subsequently processed by the DPT [47] of VGGT to generate point clouds. As shown in Fig. 4, the reconstructed mentaling point clouds roughly depict the underlying scene, where the clearer regions are typically correlated with prompt-relevant objects. This observation indicates that the 3D latents effectively encode the mental scene guided by the prompt intent. After reasoning with 3D mentaling, all three examples yield correct answers. Additional visualizations and analysis are provided in the Supp. Mat..
@@ 4.5 Ablation Study
Table 6: Ablation of different designs including 3D special token position (Token Pos.), projector and rewards in terms of $\textit{{3DThinker}-S1+S2}_{\textnormal{Qwen2.5-3B}}$.

### v4

@@ Abstract
Though recent advances in vision–language models (VLMs) have achieved remarkable progress across a wide range of multimodal tasks, understanding 3D spatial relationships from limited views remains a significant challenge. Previous reasoning methods typically rely on pure text (e.g., topological cognitive maps) or on 2D visual cues. However, their limited representational capacity hinders performance in specific tasks that require 3D spatial imagination. To address this limitation, we propose 3DThinker, a framework that can effectively exploit the rich geometric information embedded within images while reasoning, like humans do. Our framework is the first to enable 3D mentaling during reasoning without any 3D prior input, and it does not rely on explicitly labeled 3D data for training. Specifically, our training consists of two stages. First, we perform supervised training to align the 3D latent generated by VLM while reasoning with that of a 3D foundation model (e.g., VGGT). Then, we optimize the entire reasoning trajectory solely based on outcome signals, thereby refining the underlying 3D mentaling. Extensive experiments across multiple benchmarks show that 3DThinker consistently outperforms strong baselines and offers a new perspective toward unifying 3D representations into multimodal reasoning. Our code is available at https://github.com/zhangquanchen/3DThinker.
@@ 1 Introduction
Recent advances have attempted to enhance the spatial reasoning capabilities of VLMs [46, 41, 69, 35, 24, 40, 13]. As illustrated in Fig. 1, existing methods can be broadly divided into two categories. The first category performs reasoning with pure text [7, 40, 14, 46, 41] or 2D visual cues [17, 13, 72], whose representational capacity for complex spatial layouts is inherently limited. To mitigate this limitation, methods such as MindCube [83] train models to generate cognitive maps of 3D layouts; however, they rely on bird’s-eye-view (BEV) annotations to construct these maps. Ego3D [24] further employs external models—GroundingDINO [39] for referring expression comprehension (REC) and DepthAnythingv2 [77] for depth estimation, to automatically generate cognitive maps Yet, constrained by the performance of these models, such methods often fail on low-resolution or uncurated images. The second category incorporates auxiliary modalities as additional inputs (e.g., point clouds, camera parameters) [9, 33]. However, these settings restrict the model’s applicability in real-world scenarios where only monocular images are available. Moreover, several recent methods invoke external encoders or tool-usage to obtain prior information (e.g., encoded 3D tokens [21], depth maps [40, 13, 5]). Importantly, these techniques do not constitute an intrinsic capability of the model and introduce additional inference overhead.
@@ 2.1 Multimodal Reasoning
Large language models (LLMs) have experienced rapid development, and demonstrated strong performance across a wide range of tasks [11, 96, 88, 36, 85, 71, 16, 81]. Building on these advances, recent works have highlighted that in-context learning, including intermediate rationales, can significantly enhance the performance of LLMs [50, 65, 22, 64, 38, 97, 84]. Current reasoning methods can be categorized into three types: pure-text, visual, and latent reasoning.
Pure-text reasoning:  [10, 57, 30, 27, 54, 3] elicit textual step-by-step reasoning inspired by [25]. They typically rely on textual descriptions, which can limit the reasoning capabilities when dealing with visual evidence that cannot be adequately described using pure textual language. Visual reasoning: to solve the problem mentioned above, some methods integrate visual evidence directly into the reasoning trajectory, whether in multi-hop or continuous modes. Some intrinsic multi-hop methods [41, 52, 12, 63], first generate detailed visual cues within the model itself (e.g., bounding boxes, coordinates, or masks), and then the further reasoning is conducted based on these cues. Other extrinsic tool-usage methods [45, 55, 68, 95], enhance the ”think with image” capability by dynamically invoking external image tools. On the other hand, continuous methods like GRIT [20] and SIFThinker [13] generate continuous visual reasoning to enable iterative corrections during the single-step reasoning process. Latent reasoning: some studies have shown that incorporating intermediate hidden representations into LLMs can effectively enhance model capabilities [4, 79, 18, 56].  [26] replaces CoT tokens with continuous latent embeddings, allowing unconstrained reasoning in the latent space to tackle complex tasks. More recently, Mirage [80] and LVR [32] utilize special visual tokens alongside ordinary text during reasoning. They explore visual information within the model by implicitly supervising the generation of image latent, thereby enabling reasoning with 2D visual latent.
@@ 2.2 Spatial Understanding
Spatial understanding encompasses skills such as 3D imagination and spatial cognition, which are essential for perceiving and manipulating spatial relationships in both 2D and 3D environments [73, 89, 6, 44, 59, 87, 43, 90, 19, 82]. Recently, many efforts have been dedicated to evaluating the spatial understanding ability of VLMs [24, 83, 49, 92, 42, 93, 31]. Additionally, several methods have been proposed to enhance spatial understanding. For example,  [40, 5, 17, 46, 47] equip LLM with additional multiview, depth or point cloud inputs, essentially serving as input enhancement. Furthermore, 3DRS [28] introduces a teacher model for 3D supervision to achieve explicit spatial representation alignment; however, this method requires input that includes the 3D coordinates corresponding to each pixel. Moreover, VLM-3R [21] employs implicit 3D tokens from a pre-trained model (e.g., CUT3R [61]) to achieve spatial awareness by incorporating prior information, necessitating inference with an extensively 3D foundation model. Recently, methods like MindCube [83] and Ego3D-VLM [24] have facilitated spatial understanding by constructing textual cognitive maps.
@@ 3.1 Data Generation
Due to the fact that VLMs naturally only generate textual tokens, they require additional supervised training to learn how to produce interleaved reasoning patterns that incorporate 3D information. Therefore, we synthesize specific training corpora based on the raw question-answer data accompanied by multi-view images. Given the image set $\mathcal{I}=\{I_{1},I_{2},\ldots,I_{n}\}$, a question $Q$, and the ground truth response $R$, we employ a high-level model (i.e., GPT-4.1) $M$ to complete the reasoning chain. Specifically, we prompt the model $M$ to generate step-by-step reasoning that contains placeholders (3D special tokens), where these tokens represent imagined 3D scenes in the mind. Denote the response $o$ as:
@@ 3.2 Supervision for 3D Grounded Reasoning
The overall training objective incorporates both 3D alignment and textual losses, thereby enabling the model to seamlessly incorporate 3D imaging into its textual reasoning process. Here, $\lambda_{3D}$ and $\lambda_{text}$ serve as hyperparameters that balances coefficients.
@@ 3.3 Reinforced Spatial Mentaling
After the supervised training, the model has begun to exhibit the ability of 3D mentaling during the thinking process. To further optimize the 3D visual token in the reasoning process, we can extract the last layer hidden state of 3D special token $t_{3D}$ (i.e., <|latent_start|><|latent_pad|> ... <|latent_end|>) in each trajectory $o$ and perform optimization. Specifically, the projected features $F_{proj}^{RL}$ are computed based on Eq. 4 at step $t$, with VGGT features $F_{3D}$ serving as constraints during the RL stage. That is, the cosine similarity between the VGGT features and the projected features is calculated to serve as the reward $r_{\text{3D}}$.
@@ 4 Experiments
Training. Models marked with * are trained on our Large-Scale Curated Dataset, whereas those without the mark are trained exclusively on MindCube-only. For further details regarding the training dataset, please refer to the Supp. Mat..
Hyper-parameters. For 3DThinker trained on MindCube-only, in the stage 1, we set the MLP depth to 6, with the learning rate of 1e-4, latent size of 12, epoch of 10. In Eqn. 9, the hyper-parameters $\lambda_{3D},\lambda_{text}$ are uniformly set to 0.1 and 1, respectively. In the stage 2, we set the balancing coefficient of all three rewards to 1, with the learning rate of 1e-5, the rollout number of 8. Additional details are provided in the Supp. Mat..
@@
$\textit{{3DThinker}-S1}_{\textnormal{Qwen2.5-3B}}*$ 53.2 54.8 74.5 52.3 59.5 37.7 55.3
$\textit{{3DThinker}-S1+S2}_{\textnormal{Qwen2.5-3B}}*$ 59.1 60.2 78.4 58.2 64.7 41.9 60.4
@@
$\textit{{3DThinker}-S1}_{\textnormal{Qwen2.5-7B}}*$ 57.3 61.5 77.9 56.3 61.7 41.5 59.4
$\textit{{3DThinker}-S1+S2}_{\textnormal{Qwen2.5-7B}}*$ 63.7 68.3 81.1 63.3 68.6 43.3 64.7
@@ 4.1 Benchmarking Generalist VLMs
In this section, we comprehensively investigate different training stages in 3DThinker across various generalist VLMs. We conduct experiments on MindCube-Tiny [83] and Ego3D-Bench [24], both of which are designed to evaluate the spatial understanding ability from limited views.
@@ 4.3 Training Strategies
To further demonstrate the effectiveness of our training paradigm, we compare 3DThinker against several representative training strategies as shown in Tab. 4. Among them, Aug-CGMap-FFR-Out and Plain-CGMap-FFR-Out serve as SOTA baselines introduced in [83]. Specifically, Aug-CGMap-FFR-Out performs reasoning with the augmented cognitive map (camera-view information included), whereas Aug-CGMap-FFR-Out relies solely on plain cognitive maps without augmentation.
@@ 4.4 Visualization
We visualize the results in Fig. 4. During inference, we extract the last layer hidden states corresponding to the 3D special tokens. These 3D latents are projected into the VGGT feature space via the projector illustrated in Fig. 3. The projected features are subsequently processed by the DPT [48] of VGGT to generate point clouds. As shown in Fig. 4, the reconstructed point clouds roughly depict the underlying scene, where the clearer regions are typically correlated with prompt-relevant objects. This observation indicates that the 3D latents effectively encode the mental scene guided by the prompt intent. After reasoning with 3D mentaling, all two examples yield correct answers. Additional results are provided in the Supp. Mat..
@@ 4.5 Ablation Study
Table 6: Ablation of different designs including 3D special token position (Token Pos.), projector and rewards in terms of $\textit{{3DThinker}-S1+S2}_{\textnormal{Qwen2.5-3B}}$ on MindCube-Tiny.